### v2

@@ 3.1 Recurrence Operators
[…]
@@ 3.2 Solvable Operators
[…]
@@ 4 Defective Variables
Let $\mathcal{P}$ be a loop with recurrence operator $\mathcal{R}$ and $x,y\in\operatorname{Vars}(\mathcal{P})$. We say $x$ depends on $y$ if $y$ appears in a monomial in $\mathcal{R}[x]$ with non-zero coefficient. Moreover, $x$ depends linearly on $y$ if all monomials with non-zero coefficients in $\mathcal{R}[x]$ containing $y$ are linear. Analogously, $x$ depends non-linearly on $y$ if there is a non-linear monomial with non-zero coefficient in $\mathcal{R}[x]$ containing $y$. Furthermore, we consider the transitive closure for variable dependency. If $z$ depends on $y$ and $y$ depends on $x$, then $z$ depends on $x$ and, if in addition, one of these two dependencies is non-linear, then $z$ depends non-linearly on $x$. We otherwise say the dependency is linear.
@@
[…]
@@
[…]
@@ 7.1 Defective Variables in Probabilistic Loop Models
  while $\star$ do      $y\leftarrow 4y(1-y)$      $x\leftarrow x-y\ \{1/2\}\ x+y$   end while $\mathbb{E}(y_{n+1})$ $=4\mathbb{E}(y_{n})-4\mathbb{E}(y_{n}^{2})$ $\mathbb{E}(x_{n+1})$ $=\mathbb{E}(x_{n})$ $\mathbb{E}(x^{2}_{n+1})$ $=\frac{1}{2}\mathbb{E}((x_{n}-y_{n+1})^{2})+\frac{1}{2}\mathbb{E}((x_{n}+y_{n+1})^{2})$ $=\mathbb{E}(x^{2}_{n})+\mathbb{E}(y^{2}_{n+1})$
@@ 7.2 Synthesising Solvable Probabilistic Loops
[…]
@@
[…]
@@ 8 Applications of Unsolvable Operators towards Invariant Generation
[…]
@@
[…]
@@
Further examples of trace maps (Example 37 and Example 38 below) are constructed from similar substitution rules on strings of matrices. For Examples 3638, our work generates the cubic polynomial invariant
$x^{2}+y^{2}+z^{2}-2xyz=x_{0}^{2}+y_{0}^{2}+z_{0}^{2}-2x_{0}y_{0}z_{0},$
from the well-studied class of Fricke–Vogt invariants as well as higher-degree polynomial invariants.
@@ Acknowledgements
This research was supported by the Vienna Science and Technology Fund WWTF 10.47379/ICT19018 grant ProbInG, the European Research Council Consolidator Grant ARTIST 101002685, the Austrian Science Fund (FWF) project W1255-N23, and the SecInt Doctoral College funded by TU Wien.

### v3

@@ 3.1 Recurrence Operators
Consider the program $\mathcal{P}_{\textrm{SC}}$ in Figure 1(b). One can employ a recurrence operator $\mathcal{R}$ in order to capture the values of the program variables in the $n$th iteration. For $v\in\operatorname{Vars}(\mathcal{P}_{\textrm{SC}})$, $\mathcal{R}[v]$ is obtained by bottom-up substitution in the polynomial updates starting with $v$. As a result, we obtain the following system of recurrences:
@@ 3.2 Solvable Operators
Consider the probabilistic program $\mathcal{P}_{\textrm{MC}}$ of  [25, 26] modelling a non-linear Markov chain, where $\operatorname{Bernoulli}({p})$ refers to a Bernoulli distribution with parameter $p$. Here the updates to the program variables $x$ and $y$ occur simultaneously.
@@ 4 Defective Variables
Let $\mathcal{P}$ be a loop with recurrence operator $\mathcal{R}$ and $x,y\in\operatorname{Vars}(\mathcal{P})$. We say $x$ depends on $y$ if $y$ appears in a monomial in $\mathcal{R}[x]$ with non-zero coefficient. Moreover, $x$ depends linearly on $y$ if all monomials with non-zero coefficients in $\mathcal{R}[x]$ containing $y$ are linear. Analogously, $x$ depends non-linearly on $y$ if there is a non-linear monomial with non-zero coefficient in $\mathcal{R}[x]$ containing $y$.
Furthermore, we consider the transitive closure for variable dependency. If $z$ depends on $y$ and $y$ depends on $x$, then $z$ depends on $x$ and, if in addition, one of these two dependencies is non-linear, then $z$ depends non-linearly on $x$. We otherwise say the dependency is linear.
@@
From the recurrence equations of Example 3 for the program $\mathcal{P}_{\textrm{SC}}$ (see Figure 1(b)), one obtains the dependencies between the program variables of $\mathcal{P}_{\textrm{SC}}$: the program variable $w$ depends linearly on both $x$ and $y$, whilst $x$ and $y$ depend non-linearly on each other and on $w$. By definition, the partition into effective and defective variables is $E(\mathcal{P}_{\textrm{SC}})=\emptyset$ and $D(\mathcal{P}_{\textrm{SC}})=\{w,x,y\}$.
@@
Figure 2: The dependency graphs for $\mathcal{P}_{\textrm{SC}}$ and $\mathcal{P}_{\square}$ from Figure 1.
@@ 6 Synthesising Solvable Loops from Unsolvable Loops
  $z\leftarrow 0$
  while $\star$ do
     $z\leftarrow 1-z$
     $x\leftarrow 2x+y^{2}+z$
     $y\leftarrow 2y-y^{2}+2z$
  end while
Algorithm 2
@@
  $z\leftarrow 0$
  $\texttt{s}\leftarrow x_{0}+y_{0}$
  while $\star$ do
     $\begin{pmatrix}z\\ \texttt{s}\end{pmatrix}\leftarrow\begin{pmatrix}1-z\\ 2\texttt{s}-3z+3\end{pmatrix}$
  end while
@@ 7.1 Defective Variables in Probabilistic Loop Models
  while $\star$ do
     $y\leftarrow 4y(1-y)$
     $x\leftarrow x-y\ \{1/2\}\ x+y$
  end while
$\mathbb{E}(y_{n+1})$ $=4\mathbb{E}(y_{n})-4\mathbb{E}(y_{n}^{2})$
$\mathbb{E}(x_{n+1})$ $=\mathbb{E}(x_{n})$
$\mathbb{E}(x^{2}_{n+1})$ $=\frac{1}{2}\mathbb{E}((x_{n}-y_{n+1})^{2})+\frac{1}{2}\mathbb{E}((x_{n}+y_{n+1})^{2})$
$=\mathbb{E}(x^{2}_{n})+\mathbb{E}(y^{2}_{n+1})$
@@ 7.2 Synthesising Solvable Probabilistic Loops
Recall the program $\mathcal{P}_{\textrm{MC}}$ of Example 5. An invariant synthesised by our approach in Section 5 with degree $1$ is $\mathbb{E}(x_{n}-y_{n})=\frac{5^{n}}{6^{n}}(x_{0}-y_{0})$. Hence the solvable loop synthesised by Algorithm 2 for $\mathcal{P}_{\textrm{MC}}$ with input degree $1$ is
@@
where f is the fresh variable introduced by Algorithm 2 modelling $\mathbb{E}(x_{n}-y_{n})$. Our approach from Algorithm 2 synthesises this solvable loop from $\mathcal{P}_{\textrm{MC}}$, using less than $0.5$ second within our implementation (Section 9).
@@ 8 Applications of Unsolvable Operators towards Invariant Generation
As mentioned in Example 26, $\mathbb{E}(x_{n}-y_{n})=\frac{5^{n}}{6^{n}}(x_{0}-y_{0})$ is an invariant for the program $\mathcal{P}_{\textrm{MC}}$ introduced in Example 5. Closed-form solutions for higher order expressions are also available; for example,
@@
refers to the $d$th moment of $x(n)-y(n)$. While the work in [25] uses martingale theory to synthesise the above invariant (of degree 1), our approach automatically generates such invariants over higher-order moments (see Table 2). We note to this end that the defective variables in $\mathcal{P}_{\textrm{MC}}$ are precisely $x$ and $y$ as can be seen from their mutual non-linear interdependence. Namely, we have $D(\mathcal{P}_{\textrm{MC}})=\{x,y\}$ and $E(\mathcal{P}_{\textrm{MC}})=\{s\}$.
@@
Further examples of trace maps (Example 37 and Example 38 below) are constructed from similar substitution rules on strings of matrices.
Let
$I_{1}(x,y,z)$ $=x^{2}+y^{2}+z^{2}-2xyz,$
$I_{2}(x,y,z)$ $=4x^{2}y-2xz-y,$
$I_{3}(x,y,z)$ $=x^{2}+y^{2}+z^{2}-xyz-xy-xz-yz-x-y-z.$
Then our work generates the cubic polynomial invariants $I_{1}(x,y,z)-I_{1}(x_{0},y_{0},z_{0})=0$, $I_{2}(x,y,z)-I_{2}(x_{0},y_{0},z_{0})=0$, and $I_{3}(x,y,z)-I_{3}(x_{0},y_{0},z_{0})=0$ for Example 36, Example 37, and Example 38, respectively.
@@ Acknowledgements
This research was supported by the Vienna Science and Technology Fund WWTF 10.47379/ICT19018 grant ProbInG, the European Research Council Consolidator Grant ARTIST 101002685, the Austrian Science Fund (FWF) project W1255-N23, and the SecInt Doctoral College funded by TU Wien. This research was also funded in part by the Austrian Science Fund (FWF) [10.55776/F85, 10.55776/W1255], for open access purposes; the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. We thank the authors of [45] for alerting us to an error in an early version of our manuscript and the anonymous reviewers for their constructive feedback.